# Magic phase transition and non-local complexity in generalized $W$ State

### v1

@@ Abstract
Entanglement has played an important role in impetuously developing our understanding of quantum many-body systems [1, 2]. However, over the years it has become increasingly clear that entanglement alone is not able to capture every feature that differentiate quantum from classical systems[3, 4]. The most relevant example of this is the fact that entanglement alone does not guarantees the so-called Quantum Supremacy [5]. Indeed, several highly entangled quantum states can be obtained from a fully factorized state by circuits made of Clifford gates [6, 7], i.e. a series of operations that can be efficiently simulated on a classical computer. Quantum advantage is attained at the price of non Clifford resources and exponential increment in the difficulty of simulating a quantum circuit on classical computers [8, 5].
@@
The resource beyond Clifford operations is colloquially known as magic [10, 9] and its quantification is a formidable challenge for quantum information science [10]. Recently, a computationally tractable family of measures of Magic for qubit systems called the Stabilizer Rényi Entropies (SREs) [11] has been proposed. Indeed, it was proven that SREs with Rényi index greater or equal to 2 are monotones of magic for pure states [12]. Among the others, the one with Rényi index equal to 2 has acquired a prominent role since in some cases is experimentally achievable [13, 15, 14].
This family of stabilizer entropies has provided a way for analyzing the complexity of quantum states in quantum many-body systems [22, 23, 24]. In gapped local systems, it was found that the SRE of ground states follows a volume law in which the slope can be determined using single-spin expectation values [23]. On the contrary, this local behavior of the SRE disappears in the presence of long-range correlations that can be induced either by placing the system near a phase transition or, as emphasized in recent years, inducing topological frustration in the system [25, 26, 27, 28]. SRE, and more generally, other magic measures, have thus proved to be useful in the analysis of quantum many-body phases[9].
In this work, we show that one can employ SRE to characterize a mirror symmetry breaking transition in the anisotropic Heisenberg model that has so far been particularly elusive and hard to detect since, up to today, there are no known non-vanishing order parameters in the thermodynamic limit and where even the analysis of entanglement properties has not been helpful.
@@
The Hamiltonian of the 1D fully anisotropic Heisenberg chain (also known as the XYZ chain) with a global magnetic field along the $z$-axis reads
@@
Here and in the following we will always assume that $J_{x}=1$ and $|J_{y}|,\,|J_{z}|<J_{x}$. For certain values of the Hamiltonian parameter this model is exactly solvable, but we will consider it in generality and thus approach it numerically. Within Frustrated Boundary Conditions (FBCs) [44, 45, 46, 47, 48, 49], the number of sites is taken to be odd ($L=2M+1$ for $M\in\mathbb{N}$) and periodic boundary conditions are assumed ($\sigma^{\alpha}_{n}=\sigma^{\alpha}_{n+L},\ \forall n$). FBCs induce a frustration of topological origin in these spin chains, with several interesting consequences on their low-energy properties such as the introduction of an unusual long-range behavior of correlations [48] and the substitution of a gapped system with one in which the ground state is part of band of $2L$ states.
According to [31], the frustrated phase of the XYZ model divides into two regions. In the first, which includes the Ising chain, the ground state is unique and has zero momentum. In [32] it was argued that the line dividing the two regions is a second order boundary QPT, but that analysis did not account for the lattice momentum quantization. In fact, the analysis performed in [31] showed no discontinuities in the free energy and did not find any observable that survive the thermodynamic limit. In this work, we analyze such a transition, considering it as an example of a general kind of phase transitions that can be characterized in terms of a discontinuity in the SRE. To the best of our knowledge, this is the first instance of a “pure magic transition” in a deterministic quantum system, even if, a shift between local and non-local magic has been recently observed in random quantum circuits [14, 33, 34, 35].
The generality of this result comes from the unique behavior of SRE in a family of generalized $W$-states in which each element is defined by its phase. Unlike bipartite entanglement, in the thermodynamic limit, the value of magic remains a function of the phase. As we will show in the following, these states can be mapped into the elements of the lower energy band of the topologically frustrated XYZ model close to the classical point, i.e. when the Hamiltonian is close to the one of classical Ising. Therefore, a crossover in the elements of the lowest energy band is associate to a change for the SRE of the ground state. To extend this result outside the perturbative regime, we generalize the results in [24]. We provide numerical evidence that, in the thermodynamic limit, it is possible to write the SRE of the ground states of a topologically frustrated spin chain as the sum of the SRE of the ground state of the corresponding non-frustrated model plus the SRE of the associated generalized $W$-states. Since without frustration the SRE is continuous within a given phase, the SRE of the topologically frustrated chain shows the same signature of phase transition inherited from the generalized $W$-states even outside the perturbative regime.
We start our analysis by calling to mind the $W$-states [36] that play a pivotal role in quantum information [37, 38, 39] finding applications in various quantum protocols, such as anonymous transmission in quantum networks [40], quantum communication [41], and error detection [42]. The family of generalized $W$-states (gWs) that we consider in this letter reads
Here $L$ is the number of qubits in our system, $p$ is a generic phase, and $\ket{\pm}$ are the eigenstates of the Pauli operator $\sigma^{x}$ corresponding to eigenvalues ±$\pm 1$. The original $W$-states are recovered by setting $p=0$. Being a generalization, some properties of the $W$-states extend to the whole family, while others will depend on $p$. Among the properties of the $W$-states that extend to every gWs there is the entanglement, which is independent of $p$. Indeed, for any bipartition, there are only two non-vanishing eigenvalues of the reduced density matrix and are equal to $(1\pm x)/2$, where $x$ is the difference between the dimension of the two subsystems normalized by $L$.
For a pure state |$\ket{\psi}$ defined in a system of $L$ qubits, the SRE (of index $2$) is defined as
$\mathcal{M}_{2}(\ket{\psi})=-\log_{2}{\left(\frac{1}{2^{L}}\sum_{\mathcal{P}}\bra{\psi}\mathcal{P}\ket{\psi}^{4}\right)}.$ (3)
Here $\mathcal{P}=\bigotimes_{i=1}^{L}P_{i}$ runs over all possible Pauli strings, $P_{i}\in\{\sigma^{0}_{i},\sigma_{i}^{x},\sigma_{i}^{y},\sigma_{i}^{z}\}$ with $\sigma_{i}^{0}$ represents the identity operator. Although eq. (18) implies a summation of $4^{L}$ expectation values, it allows for an efficient treatment with tensor networks [17, 18, 16, 19, 20, 21].
The expression of the SRE for gWs, Eq.2 is a function on $L$ and $p$ and we evaluated it analytically to be (see Supplementary Material)
In the limit $p\to 0$ we reach the minimum of eq. (23) and recover the SRE for the $W$ state:
$\mathcal{M}_{2}(0,L)=3\log_{2}{(L)}-\log_{2}{(7L-6)}.$ (5)
Moreover, if we require that the states in (2) must be also eigenstates of the translation operator [43] then $p$ becomes the quantized momentum $(p=\frac{2\pi}{L}\ell)$ with $\ell$ integer running from 0 to $L-1$. Within this hypothesis, for $\ell\neq 0$ we have that eq. (23) reduces to
Note that this expression is independent from $\ell$, as long as it is finite: the difference $\Delta\mathcal{M}_{2}(L)=\mathcal{M}_{2}\left(\frac{2\pi}{L}\ell,L\right)-\mathcal{M}_{2}(0,L)$, represents the gap in the SRE, which is associated with the state acquiring non-zero momentum that reduces to $\log_{2}(7/6)$ in the thermodynamic limit.
The family of states in (2) plays a central role in the study of topologically frustrated 1D systems. Indeed, applying on them the (magic preserving) Clifford circuit
introduced in Ref. [24] it is possible to obtain the elements of the low energy band of the 1D topologically frustrated XYZ model close to the classical point. Here $\text{H}(j)\equiv\frac{1}{\sqrt{2}}(\sigma^{x}_{j}+\sigma^{z}_{j})$ is the Hadamard gate on the $j$-th qubit, while $\text{C}(j,l)\equiv\exp\left[\frac{\pi}{4}(1-\sigma^{x}_{j})(1-\sigma^{z}_{l})\right]$ is the CNOT gate on the $l$-th qubit controlled by the value of the $j$-th one and $\Pi^{z}=\bigotimes_{j=1}^{L}\sigma^{z}_{j}$ is the parity operator along z.
@@
Figure 1: Value of $h^{*}$ as function of $J_{y}$ and $J_{z}$ ($J_{x}$ is assumed to be equal to 1) for the Hamiltonian in eq. (1). The data is obtained numerically looking at the momentum of the ground state for a system made of $L=15$ spins. For $h^{*}>0$, choosing $|h|<h^{*}$ the ground state manifold has dimension equal to 2 and is spanned by states with finite, opposite momenta $p\neq 0$.
@@
The non-trivial response of AFM spin chains to FBC is also witnessed by an excess of bipartite entanglement beyond the area-law contribution [47, 51, 50]. While these properties characterize the whole frustrated phase, accordingly with [31], assuming $J_{z}\geq-J_{y}$ there exists a critical value of the external magnetic field $h^{*}>0$ such that for $|h|<h^{*}$ the ground state manifold becomes twofold degenerate and spanned by states with finite, opposite momenta $p\neq 0$. Interestingly, the physics of the whole frustrated phase can be described in a quasi-particle picture through a single delocalized excitation in the ground state of frustrated chains. While in most cases this excitation carries zero momentum, below $h^{*}$, where the ground-state manifold is at least two-fold degenerate, it owns a non-vanishing one.
@@
In the spirit of adiabatic continuation, let us analyze a particularly simple case where the calculations can be carried out analytically, and subsequently, we will show how the results thus obtained generalize to the entire phase. Therefore, let us focus on the case in which the system is close to the classical point. We define classical point, the case in which $J_{x}$ is the only non-vanishing Hamiltonian parameter. Indeed, in this case, eq. (1) reduces to a sum of mutually commuting terms, i.e. to a classical Hamiltonian. In this region, exploiting perturbation theory, we obtain that the elements of the lowest energy band can be written as kink states
@@
where $p$ is the quantized momentum, i.e. $p=2\pi\ell/L$, with $\ell=0,\ldots,L-1$. The kinks are embedded in Neél order states and are made of the union of two extensive sets of states defined as |$\ket{k}={\cal T}^{k}\bigotimes_{j=1}^{M}\sigma_{2j}^{z}\ket{-}^{\otimes L}$ and |$\ket{k^{\prime}}={\cal T}^{k}\bigotimes_{j=1}^{M}\sigma_{2j}^{z}\ket{+}^{\otimes L}$ with $k$ and $k^{\prime}$ running from $1$ to $L$. Turning on $J_{y}$ $J_{z}$ and/or $h$ in the proximity of the classical point we can have either a unique or a two-fold degenerate ground state. In the first case, the lowest energy is obtained by setting $p=0$, while in the other case, the two ground states display equal but opposite momenta $p$, dependent on the Hamiltonian parameters.
Therefore, it becomes crucial to have a physical quantity able to discriminate between states that yield results dependent on $p$, or at least whether, $p$ is zero or not. In [31] the authors showed how chirality, generally sensitive to the momentum, cannot be used in this case since, even if different from 00 for finite $L$, it vanishes in the thermodynamic limit. A natural second candidate is represented by the entanglement, which has often been used to reveal the presence of non-local correlations. Summarizing the results reported in the supplementary materials, we have that, limiting ourselves to partitions ($A|B$) composed of connected subsets, regardless of their dimensions, the reduced density matrix $\rho_{A}(p)=\mathrm{Tr}_{B}(\ket{\omega_{p}}\!\!\bra{\omega_{p}})$ admits only 4 non-zero eigenvalues:
$\lambda_{1,\ldots,4}$ $=$ $\frac{1}{4L}\Big{(}L+2\gamma\cos(p\chi)\pm\Big{.}$
$\pm$ .$\Big{.}\sqrt{(L-2a)^{2}+4L(1+\gamma\cos(p\chi))-4\sin^{2}(p\chi)}\Big{)}\,,$
where $a=\mathrm{dim}(A)$, $\chi=L-a$ and $\gamma=\pm 1$. From eq. (16) it is evident that the momentum dependent contributions scale at most with the inverse square-root of $L$ and thus vanish in the thermodynamic limit. Therefore, for large $L$ the entanglement does not depend on $p$. To provide an example, setting $a=(L-1)/2$ and evaluating the 2-Rényi entropy of |$\ket{\omega_{p}}$ we obtain
@@
On the contrary, magic works perfectly to detect the finite momentum. Since the states |$\ket{\omega_{p}}$ can be obtained from the |$\ket{W_{p}}$ via a Clifford circuit, they share the same value of magic. When the ground state is unique and carries zero momentum, the value of the SRE is given by eq. (5) but when the ground state acquires a finite momentum it increases by a quantity that stays finite even in the thermodynamic limit.
@@
that clearly approaches unity as $L\to\infty$ hence proving eq. (11). The data from this analysis would already be sufficient to prove that the phase transition associated with the violation of the mirror symmetry is highlighted by a finite gap of the magic. However, it is also interesting to provide direct verification. Therefore we performed a finite-size scaling analysis of the jump in magic at the transition point $h^{*}$ for several sets of parameters. To realize this analysis, we fix the values of the anisotropies, determine numerically $h^{*}$ and plotted the difference in the SRE soon after and soon before this point. In all analyzed cases, the numerical data show a power-law convergence of the amplitude of the discontinuity to the analytically computed value of $\log_{2}(7/6)$, as shown in Fig. 3. On the contrary, the discontinuity in the bipartite entanglement shows a power-law convergence to 0, implying that in the thermodynamic limit, it is unable to highlight the presence of the phase transition. This confirms that the SRE witnesses the quantum phase transition associated to the violation of the mirror symmetry in topologically frustrated spin chains, which could therefore be classified as a first-order magic (SRE) transition.
@@
Appendix A Magic phase transition and non-local complexity in generalized W State – Supplementary Meterials
A.1 Entanglement Entropies for the $\omega_{p}$ states for a generic bipartition made of convex subsystems
@@ Appendix A Magic phase transition and non-local complexity in generalized W State – Supplementary Meterials
In (A.1) $\ket{\pm}$ denote the eigenstates of $\sigma^{x}$ associated respectively to the positive/negative eigenvalue in the $x$-direction, $M=(L-1)/2$, while $T$ stands for the translation operator that shifts the state of the system by one single site towards the right. For $k=1$ the ferromagnetic defect is placed between sites $1$ and $L$ while with $k>1$ the translation operator moves it around the whole chain.
@@
$\rho_{A}=\mathrm{Tr}_{B}(\ket{\omega_{p}}\!\!\bra{\omega_{p}})=\frac{1}{2L}\left(\begin{array}[]{cccc}\mathbf{Q}^{(b)}&\mathbf{0}^{(b)}&V^{(b)}&W^{(b)}\\ \mathbf{0}^{(b)}&\mathbf{Q}^{(b)}&W^{(b)}&V^{(b)}\\ (V^{(b)})^{\dagger}&(W^{(b)})^{\dagger}&L-b&2\cos(\chi p)\\ (W^{(b)})^{\dagger}&(V^{(b)})^{\dagger}&2\cos(\chi p)&L-b\end{array}\right)$ (15)
In eq. (15) $\rho_{A}$ is a $2a\times 2a$ square matrix, and we defined $b=a-1$ and $\chi=L-a$. The reduced density matrix $\rho_{A}$ is not block diagonal but has a block structure and each one of these blocks has a quite regular structure. To begin, the matrices $\mathbf{0}^{(b)}$ and $\mathbf{Q}^{(b)}$ are both $b\times b$ square matrices. All the elements of the first are zeros, i.e. $\mathbf{0}^{(b)}_{m,n}=0\,\forall m,\,n$, while the elements of $\mathbf{Q}^{(b)}$ obey to the following law $\mathbf{Q}^{(b)}_{m,n}=\exp(-\imath(m-n)p)$. On the contrary, both $V^{(b)}$ and $W^{(b)}$ are column vectors made of $b$ rows and one single column. The $n$-th element of $V^{(b)}$ can be written as $V^{(b)}_{n}=\exp(-\imath(b+1-n)p)$, while for $W^{(b)}$ we have $W^{(b)}_{n}=\exp(-\imath(L-n)p)$

### v2

@@ Abstract
Entanglement has played an important role in impetuously developing our understanding of quantum many-body systems [1, 2]. However, over the years, it has become increasingly clear that entanglement alone is not able to capture every feature that differentiate quantum from classical systems [3, 4]. The most relevant example of this is the fact that entanglement alone does not guarantee the so-called Quantum Supremacy [5]. Indeed, several highly entangled quantum states can be obtained from a fully factorized state by circuits made of Clifford gates [6, 7], i.e. a series of operations that can be efficiently simulated on a classical computer. Quantum advantage is attained at the price of non Clifford resources and exponential increment in the difficulty of simulating a quantum circuit on classical computers [8, 5].
@@
The resource beyond Clifford operations is colloquially known as magic [10, 9] and its quantification is a formidable challenge for quantum information science [10]. Recently, a computationally tractable family of measures of Magic for qubit systems called the Stabilizer Rényi Entropies (SREs) [11] has been proposed. For a pure state |$\ket{\psi}$ defined in a system of $L$ qubits, the SRE (of order $\alpha$) is defined as
$\mathcal{M}_{2}(\ket{\psi})=\frac{1}{1-\alpha}\log_{2}{\left(\frac{1}{2^{L}}\sum_{\mathcal{P}}|\bra{\psi}\mathcal{P}\ket{\psi}|^{2\alpha}\right)}.$ (1)
Here $\mathcal{P}=\bigotimes_{i=1}^{L}P_{i}$ runs over all possible Pauli strings, $P_{i}\in\{\sigma^{0}_{i},\sigma_{i}^{x},\sigma_{i}^{y},\sigma_{i}^{z}\}$ with $\sigma_{i}^{0}$ represents the identity operator. It was proven that SREs with Rényi index greater than or equal to 2 are monotones of magic for pure states [12]. Among the others, the one with $\alpha=2$ has acquired a prominent role, since in some cases it is experimentally achievable [13, 15, 14]. Moreover, although eq. (1) implies a summation of $4^{L}$ expectation values, it allows for an efficient treatment with tensor networks [17, 18, 16, 19, 20, 21].
The stabilizer entropies provide a way for analyzing the complexity of quantum states in quantum many-body systems [22, 23, 24]. In gapped local systems, it was found that the SRE of ground states follows a volume law in which the slope can be determined using single-spin expectation values [23]. On the contrary, this local behavior of the SRE disappears in the presence of long-range correlations that can induce topological frustration in the system [25, 26, 27, 28]. In this work, we make a crucial step further, showing how it is possible to exploit the SRE to signal a phase transition associated with a mirror symmetry breaking that can be found in the topologically frustrated (TF) 1D anisotropic Heisenberg model (also known as TF XYZ model). TF models are characterized by a delocalized excitation, and such a transition is associated with the acquisition of a non-zero lattice momentum by the excitation [29]. Being associated with a single delocalized excitation, any local or string order parameter will at most show a mesoscopic behavior that vanishes in the thermodynamic limit [26, 29]. In [30], it was argued that such a transition is a second-order boundary QPT, but that analysis did not account for the quantization of momentum. Moreover, we will show that even entanglement fails to highlight the development of ground states that violate the mirror symmetry. Therefore, to the best of our knowledge, this is the first instance of a “pure magic transition” in a deterministic quantum system at equilibrium, even if a shift between local and non-local magic has been recently observed in the dynamics of random quantum circuits [14, 31, 32, 33].
The reason for this ability of the SRE is that it sums the expectation values of all possible Pauli strings. Therefore, even if individual putative order parameters display only a mesoscopic behavior, the SRE can signal the emergence of a ground state with a non-zero lattice momentum even in the thermodynamic limit. Following the same reasoning, it is conceivable that other combinations of strings, with different physical meanings, could detect this transition. The most natural example that comes to mind for our case is the momentum operator: like the SRE, when written in terms of spin operators, it is a non-local operator that is written as the sum of a super-extensive number of Pauli strings [35]. While it would detect the TF transition in a disorder-free system, the momentum operator is not clearly definable without translational invariance, while SREs are. Hence, thanks to its generality, we pose that the analysis of SREs can constitute a new, general tool for the study of unusual phase transitions characterized by non-local behaviors in complex systems defined in terms of qubits or qudits [36], in an alternative and complementary way to entanglement.
Thus, our results go well beyond TF systems, extending, for instance, to all those models characterized by a finite number of excitations, such as the ANNNI model [34]. In our case, to better understand the SRE, we exploit a mapping between the lowest energy states of the TF XYZ chain close to the classical point (where the Hamiltonian consists of mutually commuting terms) and generalized $W$-states in which a phase characterizes each element. As we will show, in the thermodynamic limit, the SRE remains phase-dependent, while the entanglement and expectation values of any Pauli string are insensitive to it.
The implications of these results are two-fold: on one side, it shows that inhomogeneous $W$-states may have higher complexity (as measured by SRE) while retaining the same entanglement properties. On the other, generalizing the results in [24] for magic and in [25] and [27] for entanglement, we will show that for the ground state of the TF XYZ spin chains, both quantities can be written as the sum as the contribution from the corresponding non-frusterated chain and that of the generalized $W$-state. In this way, the jump in complexity in the latter due to a finite phase equals that of the frustrated chain, providing a further link for this elusive quantum phase transition.
State Example Ent. SRE
|$\ket{W_{0}}$ $\begin{array}[]{lcl}\frac{1}{\sqrt{L}}&\Big{(}&\ket{10000\ldots}\\ &+&\ket{01000\ldots}\\ &+&\ket{00100\ldots}+\ldots\Big{)}\end{array}$ Eq. (3) Eq. (5)
|$\ket{W_{p}}$ $\begin{array}[]{lcl}\frac{1}{\sqrt{L}}&\Big{(}&e^{\imath p}\ket{10000\ldots}\\ &+&e^{\imath 2p}\ket{01000\ldots}\\ &+&e^{\imath 3p}\ket{00100\ldots}+\ldots\Big{)}\end{array}$ Eq. (3) Eq. (6)
|$\ket{\omega_{0}}$ $\begin{array}[]{lcl}\frac{1}{\sqrt{2L}}&\Big{(}&\ket{\uparrow\uparrow\downarrow\uparrow\downarrow\ldots}\\ &+&\ket{\downarrow\downarrow\uparrow\downarrow\uparrow\ldots}\\ &+&\ket{\downarrow\uparrow\uparrow\downarrow\uparrow\ldots}+\ldots\Big{)}\end{array}$ Eq. (10) Eq. (5)
|$\ket{\omega_{p}}$ $\begin{array}[]{lcl}\frac{1}{\sqrt{2L}}&\Big{(}&e^{\imath p}\ket{\uparrow\uparrow\downarrow\uparrow\downarrow\ldots}\\ &+&e^{\imath p}\ket{\downarrow\downarrow\uparrow\downarrow\uparrow\ldots}\\ &+&e^{\imath 2p}\ket{\downarrow\uparrow\uparrow\downarrow\uparrow\ldots}+\ldots\Big{)}\end{array}$ Eq. (10) Eq. (6)
Table 1: Recap of the states used in this work. |$\ket{W_{0}}$ and |$\ket{W_{p}}$ are respectively the standard and generalized $W$-states defined in eq. (2): they have the same amount of entanglement in the thermodynamic limit, but a different SRE for finite/zero $p$. |$\ket{\omega_{0}}$ and |$\ket{\omega_{p}}$ are linear superposition of kink states as in eq. (8) and can be obtained from their corresponding $W$-states through the Clifford circuit in eq. (11). Therefore, they share the same amount of magic, but not of entanglement. The ground state of the XYZ chain in eq. (7) has entanglement and magic amount of resource given by its non-frustrated (non-universal) counterpart plus the contribution form the corresponding |$\ket{\omega_{0,p}}$, see eq. (12).
As it is well known, $W$-states [37] play a pivotal role in quantum information [38, 39, 40] finding applications in various quantum algorithms, such as Grover’s [41] or quantum error correction [42, 43], quantum communication [44], and error detection [45]. The family of generalized $W$-states (gWs) that we consider in this paper is the one that still preserves the invariance under spatial translation that reads
Here $L$ is the number of qubits in our system, $p=2\pi\ell/L$ with integer $\ell$ running from $-\frac{L-1}{2}$ to $\frac{L-1}{2}$ given the quantized phase, and $\ket{\pm}$ are the eigenstates of the Pauli operator $\sigma^{x}$ corresponding to eigenvalues ±$\pm 1$. It is easy to see that the entanglement properties of the states in eq. (2) are independent of $p$. Indeed, for any bipartition, the reduced density matrix admits only two non-vanishing eigenvalues that coincide with the probabilities that the single excitation, represented by the state $\ket{+}$, is in one or the other subset. Such probabilities are given by $(1\pm x)/2$, where $x$ is the difference in the number of qubits between the two subsets normalized by $L$. Thus the 2-Rényi entropy at half chains is
$S_{2}(W_{p})=-\log_{2}\left[\frac{4L^{2}+(L-1)^{2}-4(L-1)L}{4L^{2}}\right]$ (3)
On the contrary, the expression of the SRE for gWs is a function of $L$ and $p$, and we evaluated it analytically to be (see appendix)
In the limit $p\to 0$ we reach the minimum of eq. (4) and recover the SRE for the homogeneous $W$-state:
$\mathcal{M}_{2}(0,L)=3\log_{2}{(L)}-\log_{2}{(7L-6)}.$ (5)
Conversely, for $\ell\neq 0$ we have that eq. (4) becomes
Note that this expression is independent of $\ell$, as long as the latter is finite: the difference $\Delta\mathcal{M}_{2}(L)=\mathcal{M}_{2}\left(\frac{2\pi}{L}\ell,L\right)-\mathcal{M}_{2}(0,L)$, represents a jump in SRE, associated to the state acquiring a non-zero phase (i.e. momentum), which reduces to $\log_{2}(7/6)$ in the thermodynamic limit.
@@
Let us now introduce our many-body system, the Hamiltonian of the 1D XYZ chain with a global magnetic field along the $z$-axis reads
@@
Here and in the following, we will always assume that $J_{x}=1$ while $|J_{y}|,\,|J_{z}|<1$. Within Frustrated Boundary Conditions (FBCs) [46, 47, 48, 25, 26, 35], the number of sites is taken to be odd ($L=2M+1$ for $M\in\mathbb{N}$) and periodic boundary conditions are assumed ($\sigma^{\alpha}_{n}=\sigma^{\alpha}_{n+L},\ \forall n$). FBCs induce a frustration of topological origin in these spin chains, with several interesting consequences on their low-energy properties, such as the introduction of unusual long-range behavior of correlations [26] and the substitution of a gapped system with one in which the ground state is part of a band of $2L$ states [28].
@@
Figure 1: Value of $h^{*}$ as function of $J_{y}$ and $J_{z}$ ($J_{x}$ is assumed to be equal to 1) for the Hamiltonian in eq. (7). The data is obtained numerically looking at the momentum of the ground state for a system made of $L=15$ spins. For $h^{*}>0$, choosing $|h|<h^{*}$ the ground state manifold has a dimension equal to 2 and is spanned by states with finite, opposite momenta $p\neq 0$.
@@
The non-trivial response of AFM spin chains to FBC is also witnessed by an excess of bipartite entanglement beyond the area-law contribution [25, 49, 27]. While these properties characterize the whole frustrated phase, accordingly with [29], assuming $J_{z}\geq-J_{y}$ there exists a critical value of the external magnetic field $h^{*}>0$ (see Fig. 1) such that for $|h|<h^{*}$ the ground state manifold becomes twofold degenerate and spanned by states with finite, opposite momenta $p\neq 0$. Interestingly, the physics of the whole frustrated phase can be described in a quasi-particle picture through a single delocalized excitation in the ground state of frustrated chains. While in most cases this excitation carries zero momentum, below $h^{*}$, where the ground-state manifold is at least two-fold degenerate, it owns a nonvanishing one.
@@
In the beginning, let us focus on a simple case where the system can be studied analytically. This case is represented by the system close to the classical point. The classical point is when $J_{x}$ is the only non-vanishing Hamiltonian parameter and eq. (7) reduces to a sum of mutually commuting terms, i.e., to a classical Hamiltonian. Close to this point, exploiting perturbation theory, we obtain that the elements of the lowest-energy band can be written as kink states
@@
where $p$ is the quantized momentum, i.e. $p=2\pi\ell/L$, with $\ell=-\frac{L-1}{2},\ldots,\frac{L-1}{2}$. The kinks are embedded in Neél order states and are made of the union of two extensive sets of states defined as |$\ket{k}={\cal T}^{k}\bigotimes_{j=1}^{M}\sigma_{2j}^{z}\ket{-}^{\otimes L}$ and |$\ket{k^{\prime}}={\cal T}^{k}\bigotimes_{j=1}^{M}\sigma_{2j}^{z}\ket{+}^{\otimes L}$ with $k$ and $k^{\prime}$ running from $1$ to $L$. Turning on $J_{y}$ $J_{z}$ and/or $h$ in the proximity of the classical point, we can have either a unique or a two-fold degenerate ground state. In the first case, the lowest energy is obtained by setting $p=0$, while in the other case, the two ground states display equal but opposite momenta $p$. Note that a $p=0$ state is invariant under the reflection of the system with respect to any of its sites, while a state with finite momentum breaks this mirror symmetry.
As shown in the appendix, limiting ourselves to partitions ($A|B$) composed of connected subsets, regardless of their dimensions, the reduced density matrix $\rho_{A}(p)=\mathrm{Tr}_{B}(\ket{\omega_{p}}\!\!\bra{\omega_{p}})$ admits only 4 non-zero eigenvalues:
$\lambda_{1,\ldots,4}$ $=$ $\frac{1}{4L}\Big{(}L+2\gamma\cos(pa)\pm\Big{.}$
$\pm$ .$\Big{.}\sqrt{(L-2a)^{2}+4L(1+\gamma\cos(pa))-4\sin^{2}(pa)}\Big{)}\,,$
where $a=\mathrm{dim}(A)$ and $\gamma=\pm 1$. From eq. (Magic phase transition and non-local complexity in generalized $W$ State) it is evident that the momentum-dependent terms scale at most with the inverse square root of $L$ and thus vanish in the thermodynamic limit. Therefore, for diverging $L$, the entanglement does not depend on $p$. To provide an example, setting $a=(L-1)/2$ and evaluating the 2-Rényi entropy of |$\ket{\omega_{p}}$ we obtain
@@
On the contrary, SRE works perfectly to detect the emergence of a ground state with a finite momentum. To verify this result, it is enough to observe that the states in (8) can be obtained from the ones in (2) by applying to the second the (SRE preserving) Clifford circuit introduced in Ref. [24]
Here $\text{H}(j)\equiv\frac{1}{\sqrt{2}}(\sigma^{x}_{j}+\sigma^{z}_{j})$ is the Hadamard gate on the $j$-th qubit, while $\text{C}(j,l)\equiv\exp\left[\imath\frac{\pi}{4}(1-\sigma^{x}_{j})(1-\sigma^{z}_{l})\right]$ is the CNOT gate on the $l$-th qubit controlled by the value of the $j$-th one and $\Pi^{z}=\bigotimes_{j=1}^{L}\sigma^{z}_{j}$ is the parity operator along z. Since the states |$\ket{\omega_{p}}$ are obtained from the |$\ket{W_{p}}$ via a Clifford circuit, they share the same value of magic. Therefore, when the ground state is unique and carries zero momentum, its value of the SRE is given by eq. (5), but when the ground state acquires a finite momentum it increases by a quantity that stays finite even in the thermodynamic limit.
@@
that clearly approaches unity as $L\to\infty$ hence proving eq. (12). The data from this analysis would already be sufficient to prove that the phase transition associated with the violation of the mirror symmetry is highlighted by a finite gap of the magic. However, it is also interesting to provide direct verification. Therefore, we performed a finite-size scaling analysis of the jump in magic at the transition point $h^{*}$ for several sets of parameters. To realize this analysis, we fix the values of the anisotropies, determine numerically $h^{*}$ and plotted the difference in the SRE soon after and soon before this point. In all analyzed cases, the numerical data show a power-law convergence of the amplitude of the discontinuity to the analytically computed value of $\log_{2}(7/6)$, as shown in Fig. 3. In contrast, the discontinuity in the bipartite entanglement shows a power-law convergence to 0, implying that, in the thermodynamic limit, it is unable to highlight the presence of the phase transition. This confirms that the SRE witnesses the quantum phase transition associated with the violation of the mirror symmetry in topologically frustrated spin chains, which could therefore be classified as a magic (SRE) transition.
Let us stress that no observable that can be written as a product of Pauli operators could detect this phase transition. In fact, in the appendix we prove, for the gWs |$\ket{W_{p}}$ in eq. (2), that if the expectation value of a string depends on the phase $p$, then it is mesoscopic and vanishes in the thermodynamic limit. Since the Clifford circuit that transforms states |$\ket{W_{p}}$ into |$\ket{\omega_{p}}$, also maps a Pauli string into another Pauli string, only operators involving an extensive amount of Pauli strings can detect a finite momentum $p\neq 0$ also in the many-body model.
@@
Appendix A Entanglement Entropies for the $\omega_{p}$ states for a generic bipartition made of convex subsystems
@@
In (A) $\ket{\pm}$ denote the eigenstates of $\sigma^{x}$ associated respectively to the positive/negative eigenvalue in the $x$-direction, $M=(L-1)/2$, while $T$ stands for the translation operator that shifts the state of the system by one single site towards the right. For $k=1$ the ferromagnetic defect is placed between sites $1$ and $L$ while with $k>1$ the translation operator moves it around the whole chain.
@@
$\rho_{A}$ $=$ $\mathrm{Tr}_{B}(\ket{\omega_{p}}\!\!\bra{\omega_{p}})$ (16)
$=$ $\frac{1}{2L}\left(\begin{array}[]{cccc}\mathbf{Q}^{(b)}&\mathbf{0}^{(b)}&V^{(b)}&W^{(b)}\\ \mathbf{0}^{(b)}&\mathbf{Q}^{(b)}&W^{(b)}&V^{(b)}\\ (V^{(b)})^{\dagger}&(W^{(b)})^{\dagger}&L-b&2\cos(ap)\\ (W^{(b)})^{\dagger}&(V^{(b)})^{\dagger}&2\cos(ap)&L-b\end{array}\right)$ (21)
In eq. (16) $\rho_{A}$ is a $2a\times 2a$ square matrix, and $b=a-1$. The reduced density matrix $\rho_{A}$ is not block diagonal but has a block structure and each one of these blocks has a quite regular structure. To begin, the matrices $\mathbf{0}^{(b)}$ and $\mathbf{Q}^{(b)}$ are both $b\times b$ square matrices. All the elements of the first are zeros, i.e. $\mathbf{0}^{(b)}_{m,n}=0\,\forall m,\,n$, while the elements of $\mathbf{Q}^{(b)}$ obey to the following law $\mathbf{Q}^{(b)}_{m,n}=\exp(-\imath(m-n)p)$. On the contrary, both $V^{(b)}$ and $W^{(b)}$ are column vectors made of $b$ rows and one single column. The $n$-th element of $V^{(b)}$ can be written as $V^{(b)}_{n}=\exp(-\imath(b+1-n)p)$, while for $W^{(b)}$ we have $W^{(b)}_{n}=\exp(-\imath(L-n)p)$